# Nanchang, China Decoupling Forgery Semantics

### v1

@@ Abstract
In this paper, we propose a novel method for detecting DeepFakes, enhancing the generalization of detection through semantic decoupling. There are now multiple DeepFake forgery technologies that not only possess unique forgery semantics but may also share common forgery semantics. The unique forgery semantics and irrelevant content semantics may promote over-fitting and hamper generalization for DeepFake detectors. For our proposed method, after decoupling, the common forgery semantics could be extracted from DeepFakes, and subsequently be employed for developing the generalizability of DeepFake detectors. Also, to pursue additional generalizability, we designed an adaptive high-pass module and a two-stage training strategy to improve the independence of decoupled semantics. Evaluation on FF++, Celeb-DF, DFD, and DFDC datasets showcases our method’s excellent detection and generalization performance. Code is available at: https://anonymous.4open.science/r/DFS-GDD-0F42.
@@ 1 Introduction
Existing methods typically fall into two categories: direct classification and shallow decoupling classification [Yan et al., 2023]. However, these approaches often suffer from interference by irrelevant content semantics or overlook common and unique forgery semantics, leading to limited generalization [Lin et al., 2024], as illustrated in Figure 1. They often perform well in intra-domain detection, i.e., with testing datasets similar to the training dataset, but experience severe performance degradation in cross-domain detection, i.e., with testing datasets different from the training dataset.
@@ 3.3 Training Stage 1
To enhance the effectiveness of semantics for forensics purposes, we design a multi-scale high-frequency feature extraction module based on the adaptive high-pass filter. We also employ the Pag feature fusion method [Xu et al., 2023] to design a multi-scale high-frequency feature fusion module, which allows us to integrate high-frequency features into the RGB stream, preventing over-reliance on specific color textures. In Section 4, we conducted ablation studies to justify the effectiveness of the proposed modules. The supplementary material illustrates the network architecture of $Encoder_{1}$ and demonstrates the process for inputs $X_{1}$ and $Xh_{1}$. The entire process is represented as:
@@ 3.4 Training Stage 2
where $\rho_{3}$ and $\rho_{4}$ are trade-off hyperparameters.

### v3

@@ Abstract
In this paper, we propose a novel method for detecting DeepFakes, enhancing the generalization of detection through semantic decoupling. There are now multiple DeepFake forgery technologies that not only possess unique forgery semantics but may also share common forgery semantics. The unique forgery semantics and irrelevant content semantics may promote over-fitting and hamper generalization for DeepFake detectors. For our proposed method, after decoupling, the common forgery semantics could be extracted from DeepFakes, and subsequently be employed for developing the generalizability of DeepFake detectors. Also, to pursue additional generalizability, we designed an adaptive high-pass module and a two-stage training strategy to improve the independence of decoupled semantics. Evaluation on FF++, Celeb-DF, DFD, and DFDC datasets showcases our method’s excellent detection and generalization performance. Code is available at: https://github.com/leaffeall/DFS-GDD.
@@ 1 Introduction
Existing methods typically fall into two categories [Yan et al., 2023]: direct classification and shallow decoupling classification. However, these approaches often suffer from interference by irrelevant content semantics or overlook common and unique forgery semantics, leading to limited generalization [Lin et al., 2024], as illustrated in Figure 1. They often perform well in intra-domain detection, i.e., with testing datasets similar to the training dataset, but experience severe performance degradation in cross-domain detection, i.e., with testing datasets different from the training dataset.
@@ 3.3 Training Stage 1
To enhance the effectiveness of semantics for forensics purposes, we design a multi-scale high-frequency feature extraction module (MHFE) based on the adaptive high-pass filter. We also employ the Pag feature fusion method [Xu et al., 2023] to design a multi-scale high-frequency feature fusion module (MHFF), which allows us to integrate high-frequency features into the RGB stream, preventing over-reliance on specific color textures. In Section 4, we conducted ablation studies to justify the effectiveness of the proposed modules. The supplementary material illustrates the network architecture of $Encoder_{1}$ and demonstrates the process for inputs $X_{1}$ and $Xh_{1}$. The entire process is represented as:
@@ 3.4 Training Stage 2
where $\rho_{5}$ and $\rho_{6}$ are trade-off hyperparameters.
@@ 5 Conclusion
Acknowledgement
This work was supported in part by the National Natural Science Foundation of China under Grant 62262041, and in part by the Jiangxi Provincial Natural Science Foundation under Grant 20232BAB202011.